# SALT: A Flexible Semi-Automatic Labeling Tool for General LiDAR Point Clouds with Cross-Scene Adaptability and 4D Consistency

### v1

@@ Abstract
keywords:
Zero-shot learning, Annotation tool, LiDAR panoptic segmentation, 4D panoptic LiDAR segmentation, Data alignment
1 Introduction
As increasingly large and high-quality datasets become available, the artificial intelligence community is experiencing unprecedented growth. The assembly of large-scale text corpora has driven advancements in large language models (LLM) [12, 63, 1, 61], while extensive image datasets have accelerated the development of vision foundation models (VFM) [31, 55, 88, 89, 54]. However, the foundational models for 3D LiDAR, a critical modality for autonomous driving and robotics, lag behind in development due to the scarcity of current datasets [27, 64, 7, 70, 4, 65, 59, 29, 49, 76].
@@ 1 Introduction
This challenge largely stems from the high cost of qualified annotation. For instance, annotating 40,000 frames can require up to 1,700 hours of labor [4]. Therefore, despite the widespread availability of various types of LiDAR sensors, the scale of public datasets remains constrained by cost and quality considerations. A promising approach to alleviate the annotation burden is to utilize a presegmentation model. However, an effective presegmentation model relies on large-scale dataset training, which seems to put us in a chicken-and-egg dilemma.
@@
2 Related Work
2.1 LiDAR Point Cloud Segmentation
Given its importance in robotics and autonomous driving applications, 3D LiDAR segmentation has experienced flourishing development. Although segmentation tasks can be categorized into semantic level, instance level, panoptic level and 4D panoptic level, they can generally be divided into four types based on the input to the networks. Point-based [52, 28, 62, 53, 83, 81, 3, 19, 47, 37], projection-based [71, 82, 51, 32, 2, 74, 14, 84, 10], voxel-based [13, 20, 86, 43, 35, 40, 41, 26] and multi-modality-based [72, 75, 44, 11, 87, 26, 80]. Despite the notable success of LiDAR point cloud segmentation, the effectiveness of automatic annotation using them directly in a zero-shot manner remains unsatisfactory.
Recently, PPT [69] and PTv3 [68] advance multi-dataset synergistic training through an effective pretraining approach called Point Prompt Training. SFPNet [65] expands dataset variety from different types of LiDAR data and introduces sparse focal point modulation to handle these variations. While these recent works have laid foundational steps toward scaling up LiDAR-based segmentation, they still rely on training within specific datasets, lacking a truly universal approach similar to SAM [31]. Further progress is hindered by the small scale of existing datasets and complicated dataset consolidation due to disparities in point cloud distribution across different LiDAR types. These limitations restrict the potential for scaling up, leaving the field waiting for a breakthrough moment.
2.2 Rethinking Feature Alignment via VFM
@@
2.3 Labeler and Interactive Segmentation
@@
3 Methodology
3.1 Problem Statement and System Overview
@@
The pipeline of our system is shown in Fig. 3. We adopt a hierarchical approach. First, we perform spatiotemporal aggregation and then decompose objects and ground into two groups (Sec. 3.2). Then, we build a data alignment strategy to convert the modality (Sec. 3.3). Finally, we use the 4D-consistent prompting strategy with SAM2 [55] to obtain the presegmentation result and convert it back to point cloud for users to finetune (Sec. 3.4).
@@
3.2 Spatiotemporal Aggregation and Decomposition
Considering the strong spatiotemporal properties of 3D LiDAR data sequence, we first integrate a robust SLAM system [73, 42] to estimate inter-frame poses, accumulate point clouds, and designate the key stamps based on pose changes, which will be used in Sec. 3.3 and Sec. 3.4. While mainstream interactive segmentation and distilled VFM methods typically process single frames, we propose that, from a general labeling perspective, point cloud accumulation effectively mitigates point sparsity, reduces variations across different types of LiDAR, enhances inter-frame consistency, and improves both efficiency and labeling accuracy. We then modify Patchwork++ [36] to fit the ground (or ceilings), preparing two distinct point cloud sets for segmentation: $L_{object}$, and $L_{ground}$. Our algorithm design primarily focuses on $L_{object}$. For each frame in $L_{object}$, we construct an accumulated frame called Superframe, and voxelize it to obtain $V_{object}$. Superframe at key stamps is called Keyframe. Note that we classify trees and other similar categories into set $L_{object}$, which are typically categorized as stuff in traditional panoptic segmentation. This operation will be addressed in Sec. 4.1.2.
3.3 Data Alignment
@@
3.3.1 Self-Adaptive Domain Transformation
@@
$\begin{split}F&=\{F_{i}=\xi(D_{i})\triangleq\tfrac{|DFT(D_{i})|}{\max(|DFT(D_{i})|)},D_{i}\in D_{sample}\},\end{split}$ (1)
@@
In order to reduce the inter-class variance, we also designed discrimination loss $l_{diss}$.
$l_{diss}=\frac{1}{N}\sum_{n=1}^{N}||f_{\theta}(F_{n,cls=k})-fea_{cls=k}||_{2}^{2},$ (5)
@@
We also add reconstruction head and calculate the mean square error reconstruction loss $l_{reon}$.
$l_{reon}=\frac{1}{N}\sum_{n=1}^{N}||f_{r}(f_{\theta}(F_{n}))-\xi(x_{n})||_{2}^{2},$ (6)
@@
$loss_{stage2}$ $={\omega}_{1}l_{ce}+{\omega}_{2}l_{reon}+{\omega}_{3}l_{diss}.$ (8)
@@
1:Input: SA-V dataset, $V_{object}$
2:Output: Aligned data
3:Pretrain Stage:
4:# For each VFM, only a single training session is required.
5:Pseudo-Label generation
6:Train $f_{\theta}(\cdot)$ with Eq. (7)
7:Train $f_{\theta}(\cdot)$ with Eq. (8)
8:Transform Stage:
9:# Optimize for each sequence.
10:Initialize parameter for primary camera
11:while $t$ and $\boldsymbol{\alpha}$ not converged do
12:     Do pseudo-color step
13:     Perform DFT
14:     for all $V_{object}$ in batch do
15:         Calculate Eq. (9)
16:     end for
17:     Update $t$
18:     for all $V_{object}$ in batch do
19:         Calculate Eq. (10)
20:     end for
21:     Update $\boldsymbol{\alpha}$
22:end while
23:Apply projection with $t$ and $\boldsymbol{\alpha}$
24:Process data from other pseudo-cameras
3.3.2 Pseudo-Color Generation
@@
3.4 Zero-Shot Segmentation with 4D Consistency
Compared to SAM [31], SAM2 [55] requires a label to be specified when prompted. We adopt a frame-by-frame automatic prompting approach. Specifically, we employ DBSCAN [16] to derive bi-level cluster centers for prompt calculation at Keyframes. We match the low-density clustering centers with the high-density clustering centers, and use the high-density clustering centers as the positive prompt. For each positive prompt, its negative prompt comes from the neighbor of the low-density center it matches. The prompt points are then transformed across Superframes within the Keyframe’s neighborhood through coordinate transformations provided in Sec. 3.2 and are finally projected onto a pseudo-image as 4D-consistent prompt points. Leveraging SAM2’s refinement prompts, the memory bank mechanism, and the Keyframe’s informative and representative nature, this method enables us to achieve spatiotemporal consistent prompts across frames.
@@ 3.4 Zero-Shot Segmentation with 4D Consistency
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets and Platform
@@
4.1.2 Metrics and Evaluation
@@
[…]
Table 1: Panoptic segmentation on SemanticKITTI and nuScenes validation sets. We compared SALT with fully supervised methods and self-supervised method. Note that all results are obtained from the literature.
@@
4.2 Labeling Results
SemanticKITTI. Segmentation results for SemanticKITTI validation sets are shown in Tab. 1. Compared to the self-supervised zero-shot method SAL [48], our approach not only eliminates the need for training on this dataset’s LiDAR data but also achieves a significant 18.0% $PQ$ and 18.4% $mIoU$ improvement. Even compared to supervised methods, our approach achieves 67.6 ~ 74.2% of the performance of SOTA methods in terms of $PQ$ in LiDAR-only mode. With the inclusion of the camera modality, our annotation performance improves slightly by 0.4% in $PQ$ and 1.2% in $mIoU$. This further confirms that our data alignment approach offers a significant advantage over feature alignment methods in reducing dependence on calibrated cameras, particularly in datasets like SemanticKITTI, where camera data is limited. In particular, if the data collection location lacks lighting conditions (such as at night), existing methods relying on camera distillation will become completely unusable.
4D-consistent Segmentation and Tracking are crucial for annotation tools. As shown in Tab. 2, our method achieves 31.5% $LSTQ$, approximately half the performance of supervised approaches, with significant room for improvement in $S_{assoc}$. Further analysis indicates that the primary limitation stems from the current tracking abilities of SAM2. We anticipate that advancements in VFM will enhance our framework, leading to more effective annotation tools in the future.
nuScenes. In contrast to SemanticKITTI, nuScenes contains more calibrated cameras but more sparse LiDAR data. More camera allows the feature alignment method to perform better. However, our method still outperforms by 0.3% in $PQ$ with only LiDAR data and 3.0% with additional camera as shown in Tab. 1. Combined with the previous results in SemanticKITTI, it can be seen that as a method developed for LiDAR annotation tools, the performance of our method is more related to the quality of LiDAR data itself.
@@ 4.2 Labeling Results
[…]
Table 2: 4D Panoptic Segmentation on SemanticKITTI validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
S.MID. S.MID only provides hybrid-solid LiDAR data and corresponding semantic ground truth labels. Therefore, we only compare the semantic segmentation results in terms of $mIoU$ as shown in Tab. 3. Compared to SemanticKITTI and nuScenes, S.MID presents two key challenges. First, in industrial scenarios, objects with different semantics appear at varying heights in the same location, leading to occlusion and clustering issues. Second, the point cloud distribution of hybrid-solid LiDAR is inherently random, posing significant difficulties even for supervised methods. In such a challenging industrial scenario, our method achieves 64.1 ~ 68.2% of the performance of supervised methods, which demonstrate the robustness of our approach to different scenarios and LiDAR types.
@@
[…]
Table 3: Semantic Segmentation on S.MID validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
@@
[…]
Table 4: Semantic Segmentation on SemanticKITTI-16 validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
@@
[…]
Table 5: Panoptic Segmentation on indoor Low-Resolution LiDAR and outdoor combined LiDAR sets.
@@
[…]
Table 6: A comparison of representative LiDAR point cloud annotation systems. The symbol “-” indicates that the evaluation is not applicable. The number of “+” represents the performance. “Scope” means the scope of automatic annotation by the tool during the user’s first interaction with the object or scene. “ZSL Ability” means zero-shot ability.
@@
Table 7: Ablation study for core optimization module in data alignment.
@@
Table 8: Ablation study for automatic label merging methods.
Low-Resolution LiDAR. The existing datasets are equipped with LiDAR that has at least 32 channels. However, in practical applications, many devices are equipped with the low-cost, low-resolution LiDAR. Therefore, we also conduct experiments on SemanticKITTI-16. As shown in Tab. 4, our method achieves 28.2% $mIoU$. Although the performance of this challenging experiment decreases compared to the original SemanticKITTI, supervised methods also exhibit degraded performance on the same benchmark. Therefore, SALT still demonstrates a certain level of competitiveness on low-resolution data. Additionally, we collected 16-beam LiDAR data in an indoor environment and compared the presegmentation results with the manually annotated results. As shown in Tab. 5, under such challenging data conditions, our method achieves 38.0% $PQ$. After validating our method in autonomous driving and industrial scenarios, we further demonstrated its cross-scenario annotation capability in indoor environments.
Combined LiDAR sensors. Many platforms works in safety-critical scenarios are equipped with multiple types of LiDAR simultaneously. The varying characteristics of each LiDAR introduce anisotropic density variations in the point cloud data, posing additional challenges for adaptive annotation tools. The experimental results are shown in Tab. 5, where our method achieves 52.5% $PQ$ and 54.1% $PQ$ with camera. SALT achieves nearly 40~50% of the manual annotation quality across multiple scenarios and various LiDAR types, demonstrating its generalizability.
@@
4.3 Ablation Studies
In order to evaluate the performance of each design within our tools, we carry out two groups of ablation experiments utilizing the SemanticKITTI validation set as shown in Tab. 7 and Tab. 8.
4.3.1 Better Views Yield Closer Textures
We first validate the effectiveness of our self-supervised distance-based iterative optimization from Eq. (9) and Eq. (10) in determining a suboptimal projection during modality transformation. We choose the BEV perspective, known for its high informativeness and strong separation, and use a perspective similar to real cameras in the dataset as our baseline. Compared to SALT’s sequentially adaptive suboptimal view, the two commonly used projection views show a significant performance drop: -22.7% and -23.1% $PQ$ for panoptic segmentation task and -26.3% and -16.6% $LSTQ$ for 4D panoptic LiDAR segmentation as shown in Tab. 7. These results prove that the pseudo-image in SALT aligns more closely with the SA-V dataset distribution than other projections, ultimately leading to better segmentation and tracking performance.
4.3.2 4D NMS and Smoothing Make a Good Team

### v2

@@ Abstract
Index Terms:
Zero-shot learning, annotation tool, LiDAR panoptic segmentation, 4D panoptic LiDAR segmentation, data alignment.
I Introduction
As increasingly large and high-quality datasets become available, the artificial intelligence and robotics communities are experiencing unprecedented growth. The assembly of large-scale text corpora has driven advancements in large language models (LLM) [1, 2, 3, 4], while extensive image datasets have accelerated the development of vision foundation models (VFM) [5, 6, 7, 8, 9]. However, the foundational models for 3D LiDAR, a critical modality for autonomous driving and robotics [10], lag behind in development due to the scarcity of current datasets [11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
@@ I Introduction
This challenge largely stems from the high cost of qualified annotation [21]. For instance, annotating 40,000 frames can require up to 1,700 hours of labor [15]. Therefore, despite the widespread availability of various types of LiDAR sensors, the scale of public datasets remains constrained by cost and quality considerations. A promising approach to alleviate the annotation burden is to utilize a presegmentation model. However, an effective presegmentation model relies on large-scale dataset training, which seems to put us in a chicken-and-egg dilemma.
@@
II Related Work
II-A LiDAR Point Cloud Segmentation
Given its importance in robotics and autonomous driving applications, 3D LiDAR segmentation has experienced flourishing development. Although segmentation tasks can be categorized into semantic level, instance level, panoptic level and 4D panoptic level, they can generally be divided into four types based on the input to the networks. Point-based [30, 31, 32, 33, 34, 35, 36, 37, 38], projection-based [39, 40, 41, 42, 43, 44, 45, 46, 47, 10], voxel-based [48, 49, 50, 51, 52, 53, 54, 55] and multi-modality-based [56, 57, 58, 59, 60, 55]. Despite the notable success of LiDAR point cloud segmentation, the effectiveness of automatic annotation using them directly in a zero-shot manner remains unsatisfactory.
Recently, PPT [61] and PTv3 [62] advance multi-dataset synergistic training through an effective pretraining approach called Point Prompt Training. SFPNet [16] expands dataset variety from different types of LiDAR data and introduces sparse focal point modulation to handle these variations. COLA [21] performs multi-source domain generalization by introducing the same coarse label to multiple datasets. This pretraining method enables the network to overcome the domain differences between different types of LiDAR to a certain extent and makes the subsequent fintuning of a single dataset perform better. While these recent works have laid foundational steps toward scaling up LiDAR-based segmentation, they still rely on training within specific datasets, lacking a truly universal approach similar to SAM [5]. Further progress is hindered by the small scale of existing datasets and complicated dataset consolidation due to disparities in point cloud distribution across different LiDAR types. These limitations restrict the potential for scaling up, leaving the field waiting for a breakthrough moment.
II-B Rethinking Feature Alignment via VFM
@@
II-C Labeler and Interactive Segmentation
@@
III Method
III-A Problem Statement and System Overview
@@
The pipeline of our system is shown in Fig. 3. We adopt a hierarchical approach. First, we perform spatiotemporal aggregation and then decompose objects and ground into two groups (Sec. III-B). Then, we build a data alignment strategy to convert the modality (Sec. III-C). Finally, we use the 4D-consistent prompting strategy with SAM2 [6] to obtain the presegmentation result and convert it back to point cloud for users to finetune (Sec. III-D).
@@
III-B Spatiotemporal Aggregation and Decomposition
Considering the strong spatiotemporal properties of 3D LiDAR data sequence, we first integrate a robust SLAM system [69, 70] to estimate inter-frame poses, accumulate point clouds, and designate the key stamps based on pose changes, which will be used in Sec. III-C and Sec. III-D. While mainstream interactive segmentation and distilled VFM methods typically process single frames, we propose that, from a general labeling perspective, point cloud accumulation effectively mitigates point sparsity, reduces variations across different types of LiDAR, enhances inter-frame consistency, and improves both efficiency and labeling accuracy. We then modify Patchwork++ [71] to fit the ground (or ceilings), preparing two distinct point cloud sets for segmentation: $L_{object}$, and $L_{ground}$. Our algorithm design primarily focuses on $L_{object}$. For each frame in $L_{object}$, we construct an accumulated frame called Superframe, and voxelize it to obtain $V_{object}$. Superframe at key stamps is called Keyframe. Note that we classify trees and other similar categories into set $L_{object}$, which are typically categorized as stuff in traditional panoptic segmentation. This operation will be addressed in Sec. IV-A2.
III-C Data Alignment
@@
III-C1 Self-Adaptive Domain Transformation
@@
$\begin{split}F&=\{F_{i}=\xi(D_{i})=\tfrac{|DFT(D_{i})|}{\max(|DFT(D_{i})|)},D_{i}\in D_{sample}\},\end{split}$ (1)
@@
In order to reduce the inter-class variance, we also designed discrimination loss $l_{dis}$.
$l_{dis}=\frac{1}{N}\sum_{n=1}^{N}||f_{\theta}(F_{n,cls=k})-fea_{cls=k}||_{2}^{2},$ (5)
@@
We also add reconstruction head and calculate the mean square error reconstruction loss $l_{recon}$.
$l_{recon}=\frac{1}{N}\sum_{n=1}^{N}||f_{r}(f_{\theta}(F_{n}))-\xi(x_{n})||_{2}^{2},$ (6)
@@
$loss_{stage2}$ $={\omega}_{1}l_{ce}+{\omega}_{2}l_{recon}+{\omega}_{3}l_{dis}.$ (8)
@@
  Input: SA-V dataset, $V_{object}$
  Output: Aligned data
  Pretrain Stage:
       # For each VFM, only a single training session is required.
       Pseudo-Label generation
       Train $f_{\theta}(\cdot)$ with Eq. (7)
       Train $f_{\theta}(\cdot)$ with Eq. (8)
  Transform Stage:
       # Optimize for each sequence.
       Initialize parameter for primary camera
  while $t$ and $\boldsymbol{\alpha}$ not converged do
          Do pseudo-color step
          Perform DFT
     for all $V_{object}$ in batch do
                 Calculate Eq. (9)
     end for
          Update $t$
     for all $V_{object}$ in batch do
                 Calculate Eq. (10)
     end for
          Update $\boldsymbol{\alpha}$
  end while
  Apply projection with $t$ and $\boldsymbol{\alpha}$
  Process data from other pseudo-cameras
III-C2 Pseudo-Color Generation
@@
III-D Zero-Shot Segmentation with 4D Consistency
Compared to SAM [5], SAM2 [6] requires a label to be specified when prompted. We adopt a frame-by-frame automatic prompting approach. Specifically, we employ DBSCAN [78] to derive bi-level cluster centers for prompt calculation at Keyframes. We match the low-density clustering centers with the high-density clustering centers, and use the high-density clustering centers as the positive prompt. For each positive prompt, its negative prompt comes from the neighbor of the low-density center it matches. The prompt points are then transformed across Superframes within the Keyframe’s neighborhood through coordinate transformations provided in Sec. III-B and are finally projected onto a pseudo-image as 4D-consistent prompt points. Leveraging SAM2’s refinement prompts, the memory bank mechanism, and the Keyframe’s informative and representative nature, this method enables us to achieve spatiotemporal consistent prompts across frames.
@@ III-D Zero-Shot Segmentation with 4D Consistency
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets and Platform
@@
IV-A2 Metrics and Evaluation
TABLE I: Panoptic segmentation on SemanticKITTI and nuScenes validation sets. We compared SALT with fully supervised methods and self-supervised method. Note that all results are obtained from the literature.
@@ IV-A2 Metrics and Evaluation
IV-B Labeling Results
SemanticKITTI. Segmentation results for SemanticKITTI validation sets are shown in Tab. I. Compared to the self-supervised zero-shot method SAL [26], our approach not only eliminates the need for training on this dataset’s LiDAR data but also achieves a significant 18.0% $PQ$ and 18.4% $mIoU$ improvement. Even compared to supervised methods, our approach achieves 67.6 ~ 74.2% of the performance of SOTA methods in terms of $PQ$ in LiDAR-only mode. With the inclusion of the camera modality, our annotation performance improves slightly by 0.4% in $PQ$ and 1.2% in $mIoU$. This further confirms that our data alignment approach offers a significant advantage over feature alignment methods in reducing dependence on calibrated cameras, particularly in datasets like SemanticKITTI, where camera data is limited. In particular, if the data collection location lacks lighting conditions (such as at night), existing methods relying on camera distillation will become completely unusable.
4D-consistent Segmentation and Tracking are crucial for annotation tools. As shown in Tab. II, our method achieves 31.5% $LSTQ$, approximately half the performance of supervised approaches, with significant room for improvement in $S_{assoc}$. Further analysis indicates that the primary limitation stems from the current tracking abilities of SAM2. We anticipate that advancements in VFM will enhance our framework, leading to more effective annotation tools in the future.
nuScenes. In contrast to SemanticKITTI, nuScenes contains more calibrated cameras but more sparse LiDAR data. More camera allows the feature alignment method to perform better. However, our method still outperforms by 0.3% in $PQ$ with only LiDAR data and 3.0% with additional camera as shown in Tab. I. Combined with the previous results in SemanticKITTI, it can be seen that as a method developed for LiDAR annotation tools, the performance of our method is more related to the quality of LiDAR data itself.
TABLE II: 4D Panoptic Segmentation on SemanticKITTI validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
@@ IV-B Labeling Results
S.MID. S.MID only provides hybrid-solid LiDAR data and corresponding semantic ground truth labels. Therefore, we only compare the semantic segmentation results in terms of $mIoU$ as shown in Tab. III. Compared to SemanticKITTI and nuScenes, S.MID presents two key challenges. First, in industrial scenarios, objects with different semantics appear at varying heights in the same location, leading to occlusion and clustering issues. Second, the point cloud distribution of hybrid-solid LiDAR is inherently random, posing significant difficulties even for supervised methods. In such a challenging industrial scenario, our method achieves 64.1 ~ 68.2% of the performance of supervised methods, which demonstrate the robustness of our approach to different scenarios and LiDAR types.
TABLE III: Semantic Segmentation on S.MID validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
@@
TABLE IV: Semantic Segmentation on SemanticKITTI-16 validation set. We compared SALT with fully supervised methods. Note that all results are obtained from the literature.
@@
TABLE V: Panoptic Segmentation on indoor Low-Resolution LiDAR and outdoor combined LiDAR sets.
@@
TABLE VI: A comparison of representative LiDAR point cloud annotation systems. The symbol “-” indicates that the evaluation is not applicable. The number of “+” represents the performance. “Scope” means the scope of automatic annotation by the tool during the user’s first interaction with the object or scene. “ZSL Ability” means zero-shot ability.
@@
TABLE VII: Ablation study for core optimization module in data alignment.
@@
TABLE VIII: Ablation study for automatic label merging methods.
@@
Low-Resolution LiDAR. The existing datasets are equipped with LiDAR that has at least 32 channels. However, in practical applications, many devices are equipped with the low-cost, low-resolution LiDAR. Therefore, we also conduct experiments on SemanticKITTI-16. As shown in Tab. IV, our method achieves 28.2% $mIoU$. Although the performance of this challenging experiment decreases compared to the original SemanticKITTI, supervised methods also exhibit degraded performance on the same benchmark. Therefore, SALT still demonstrates a certain level of competitiveness on low-resolution data. Additionally, we collected 16-beam LiDAR data in an indoor environment and compared the presegmentation results with the manually annotated results. As shown in Tab. V, under such challenging data conditions, our method achieves 38.0% $PQ$. After validating our method in autonomous driving and industrial scenarios, we further demonstrated its cross-scenario annotation capability in indoor environments.
Combined LiDAR sensors. Many platforms works in safety-critical scenarios are equipped with multiple types of LiDAR simultaneously. The varying characteristics of each LiDAR introduce anisotropic density variations in the point cloud data, posing additional challenges for adaptive annotation tools. The experimental results are shown in Tab. V, where our method achieves 52.5% $PQ$ and 54.1% $PQ$ with camera. SALT achieves nearly 40~50% of the manual annotation quality across multiple scenarios and various LiDAR types, demonstrating its generalizability.
@@
IV-C Ablation Studies
In order to evaluate the performance of each design within our tools, we carry out two groups of ablation experiments utilizing the SemanticKITTI validation set as shown in Tab. VII and Tab. VIII.
IV-C1 Better Views Yield Closer Textures
We first validate the effectiveness of our self-supervised distance-based iterative optimization from Eq. (9) and Eq. (10) in determining a suboptimal projection during modality transformation. We choose the BEV perspective, known for its high informativeness and strong separation, and use a perspective similar to real cameras in the dataset as our baseline. Compared to SALT’s sequentially adaptive suboptimal view, the two commonly used projection views show a significant performance drop: -22.7% and -23.1% $PQ$ for panoptic segmentation task and -26.3% and -16.6% $LSTQ$ for 4D panoptic LiDAR segmentation as shown in Tab. VII. These results prove that the pseudo-image in SALT aligns more closely with the SA-V dataset distribution than other projections, ultimately leading to better segmentation and tracking performance.
IV-C2 4D NMS and Smoothing Make a Good Team